# Bayesian Additive Regression Trees for functional ANOVA model

### v2

@@ 5.1 Posterior Concentration Rate
For technical simplicity, we let $\mathcal{X}_{j}=[0,1]$ for all $j\in[p].$ We consider the populationally identifiable ANOVA decomposition:
$f_{0}(\mathbf{x})=\sum_{S\subseteq[p]}f_{0,S}(\mathbf{x}_{S}),$
where $f_{0,S},S\subseteq[q]$ satisfy the populational identifiability condition ($\mathbb{P}_{\mathbf{X}}$-identifiability condition). We assume the following regularity conditions.
@@
    There exists a density $p_{\mathbf{X}}$ of $\mathbb{P}_{\mathbb{X}}$ with respect to the Lebesgue measure on $\mathbb{R}^{p}$ such that $0<\inf_{\mathbb{x}\in\mathcal{X}}p_{\mathbf{X}}(\mathbb{x})\leq\sup_{\mathbb{x}\in\mathcal{X}}p_{\mathbf{X}}(\mathbb{x})<\infty.$
@@
    Additionally, we assume that $\|f_{0,S}\|_{\infty}\leq F$ for some positive constant $F$. For these assumptions, we write $f_{0,S}\in\mathcal{H}^{\alpha}_{F}.$ Furthermore, we write $f_{0}\in\mathcal{H}^{\alpha}_{0,F}$ if $f_{0,S}\in\mathcal{H}^{\alpha}_{F}$ for all $S\subseteq[p]$.
@@
    There exist $0<\sigma_{\min}^{2}<\sigma_{\max}^{2}<\infty$ such that $\sigma_{0}^{2}\in(\sigma^{2}_{\min},\sigma^{2}_{\max}).$
@@
Theorem 5.1 (Posterior concentration of ANOVA-BART).
Assume that (J.1), (J.2), (J.3) and (J.4) hold. Then, for given $\xi>\max\{2^{p}F,1/\sigma^{2}_{\min},\sigma^{2}_{\max}\},$ we have
$\pi_{\xi}\big{\{}(f,\sigma^{2}):\|f_{0}-f\|_{2,n}+|\sigma_{0}^{2}-\sigma^{2}\big{|}>B_{n}\epsilon_{n}\big{|}\mathbf{X}^{(n)},Y^{(n)}\big{\}}\xrightarrow{}0,$
for any $B_{n}\xrightarrow{}\infty$ in $\mathbb{P}_{0}^{n}$ as $n\xrightarrow{}\infty,$ where $\epsilon_{n}=n^{-\frac{\alpha}{2\alpha+p}}(\log n)^{1\over 2}.$
Note that our prior is assigned to satisfy the $\mu_{n}$-identifiability condition. However, the posterior concentrate to the true function, which satisfies the $\mathbb{P}_{\mathbf{X}}$-identifiability condition. This discrepancy constitutes one of the technically delicate points of the proofs of Theorem 5.1 and Theorem 5.2 that follows.
5.2 Comparison with BART
Posterior concentration rates of BART are derived in bforest ; onBART ; artbart . Even though ANOVA-BART is a special form of BART, the way of approximation $f_{0}$ is much different from BART. In particular, the number of decision trees increases as the sample size increases in ANOVA-BART while the sizes of each decision tree increase in BART.
ANOVA-BART approximates $f_{0}$ using a sum of many simple decision trees, whereas BART approximates $f_{0}$ using a finite number of large decision trees. That is, we have to prove that the sum of fixed size small identifiable binary-product trees can approximate a smooth function well. For this purpose, we first approximate a given function by a large decision tree as is done by bforest . Then, we prove that this large decision tree can be represented by the sum of many small identifiable binary-product trees, which is technically quite involved. See Section F of Supplementary Material.
An important practical advantage of ANOVA-BART compared to BART is Theorem 5.2 below, which derives the posterior concentration of each interaction. An obvious applications of Theorem 5.2 is to screen out unnecessary interactions a posteriori.
Theorem 5.2 (Posterior concentration of each interaction).
@@ 5.2 Comparison with BART
Then, for $S\subseteq[p]$ and $\xi>\max\{2^{p}F,1/\sigma^{2}_{\min},\sigma^{2}_{\max}\},$ we have
@@
$\pi\{\|f_{S}\|_{2,n}>\epsilon_{n}\log n|\mathcal{D}\}<\delta$
for a given positive real number $\delta\in(0,1/2).$ Theorem 5.2 implies that this post-hoc component selection procedure is selection-consistent (i.e. deleting all unnecessary components).
@@ 6 Experiments
This section presents the results of numerical experiments of ANOVA-BART. In Section 6.1, we conduct an analysis of synthetic data, while we focuse on real data analysis in Section 6.2.
We consider BART (BART ), SSANOVA (SSANOVA ), MARS (MARS ) and NAM (nam ) as the baseline methods. We use ‘BayesTree’ R package (chipman2016package ), ‘gss’ R package (gu2014smoothing ),‘earth’ R package (milborrow2017earth ) and the official code in https://github.com/AmrMKayid/nam for implementing BART, SSANOVA, MARS and NAM, respectively. For NAM, we extend the code to implement NAMs to include the second order interactions. The detailed descriptions of the selection of hyperparameters in ANOVA-BART and baseline methods are presented in Section H of Supplementary Material. Note that while SSANOVA, MARS and NAM require specifying the maximum order of the components to be estimated, ANOVA-BART does not require the predefined maximum order.
6.1 Synthetic data analysis
We evaluate the performance of ANOVA-BART in view of prediction and component selection by analyzing synthetic data. To do this, we generate data using the Friedman’s test function (MARS ; BART ; linero2018bayesian ) defined as
@@ 6.1 Synthetic data analysis
where $\epsilon\sim N(0,\sigma_{\epsilon}^{2})$. The error variance $\sigma_{\epsilon}^{2}$ is set to make the signal-to-noise ratio be 5. For training data, we generate $\mathbf{x}_{i}=(x_{i1},...,x_{ip})^{\top}$, i=1,,1000i=1,...,1000 from the uniform distribution on $(0,1)^{p}$ and then generate the output using only the first 5 covairates. This means that all covariates except the first five are noise.
@@ 6.1.1 Prediction performance
For all methods, hyperparameters are selected via 5-fold cross-validation on the training data. Independently generated test data of size 10,000 are used to compute the Root Mean Square Error (RMSE) as a measure of predictive performance. For ANOVA-BART and BART, we use the Bayes estimators.
Table 1 presents the averages and standard errors of the RMSEs divided by the RMSE of ANOVA-BART based on 5 repetitions of the simulation. The optimal results are highlighted in bold. The results amply support that ANOVA-BART is favorably compared to the baseline emthods in prediction performance.
Table 1: Averages and standard errors of normalized RMSE on the synthetic data.
ANOVA-BART BART SSANOVA MARS NAM
p=10 1.000 (—) 1.094 (0.02) 1.096 (0.03) 1.065 (0.01) 1.096 (0.03)
p=50 1.000 (—) 1.065 (0.01) 1.200 (0.02) 1.015 (0.03) 1.327 (0.03)
p=100 1.000 (—) 1.051 (0.01) 1.142 (0.02) 0.944 (0.02) 1.193 (0.01)
@@ 6.1.2 Component selection
To examine how effectively ANOVA-BART selects the true signal components, we conduct a simulation. We use the $l_{2}$ norm of each estimated component (i.e., $\|f_{S}\|_{2,n}$) as the important score, i.e., if $\|f_{S}\|_{2,n}$ large, we consider $f_{S}$ to be important. We select the top 10 components based on the normalized important scores, which are presented in Figure 2 for various values of $p.$ The red bars correspond to the signal components, while the blue bars are noisy ones. The results strongly indicate that ANOVA-BART can detect signal components very well.
@@ 6.2 Real data analysis
Real data Size Dimension of covariates
boston 506 13
abalone 4,177 10
bbb 208 134
servo 167 12
mpg 392 7
We analyze 5 real data sets including boston, abalone, bbb, servo and mpg data sets. The four data sets except bbb are analyzed in BART while bbb is examined in linero2018bayesian . Table 2 summarizes the descriptions of the 5 datasets. We split each data into 80% training data and 20% test data, and repeat this random split 5 times to obtain 5 prediction measures.
@@ 6.2.1 Prediction performance
Table 3 presents the averages and standard errors of the RMSE values divided by the RMSE of ANOVA-BART for 5 datasets, where the best results are highlighted by bold. Similar to the synthetic data, ANOVA-BART is superior in prediction for real data.
Table 3: Averages and standard errors of normalized RMSE on real datasets.
Real data ANOVA-BART BART SSANOVA MARS NAM
Boston 1.000 (—) 1.179 (0.11) 1.294 (0.10) 1.389 (0.09) 1.038 (0.05)
abalone 1.000 (—) 1.022 (0.03) 0.993 (0.03) 0.993 (0.01) 0.917 (0.04)
bbb 1.000 (—) 1.053 (0.01) 1.089 (0.02) 1.396 (0.08) 1.112 (0.03)
servo 1.000 (—) 1.082 (0.10) 2.595 (0.08) 1.393 (0.14) 2.528 (0.08)
mpg 1.000 (—) 1.266 (0.15) 1.447 (0.17) 1.413 (0.18) 1.286 (0.14)
@@ 6.2.2 Uncertainty quantification
An important advantage of Bayesian methods compared to frequentist’s counterparts is superior performance of uncertainty quantification. We compare ANOVA-BART and BART in view of uncertainty quantification. As a measure of uncertainty quantification, we consider Continuous Ranked Probability Score (CRPS, gneiting2007strictly ). For a given test sample $(\mathbb{x},y),$ CRPS is defined as
@@
where $F_{\mathbb{x}}$ is the predictive distribution of $Y$ given $\mathbb{X}=\mathbb{x}.$ Usually, we use the average of CRPS on test data. A smaller CRPS means better uncertainty quantification.
Table 4 presents the average and standard errors of 5 CRPS values normalized by the CRPS of ANOVA-BART obtained by 5 random splits of training and test data with the best results highlighted in bold. The results suggest that ANOVA-BART is better than BART in uncertainty quantification, too.
Table 4: Averages and standard errors of normalized CRPS on real datasets.
Real data ANOVA-BART BART
Boston 1.000 (—) 1.087 (0.04)
abalone 1.000 (—) 0.859 (0.04)
bbb 1.000 (—) 1.025 (0.03)
servo 1.000 (—) 1.220 (0.02)
mpg 1.000 (—) 1.006 (0.04)

### v3

@@ 5.1 Posterior Concentration Rate
For technical simplicity, we let $\mathcal{X}_{j}=[0,1]$ for all $j\in[p].$ We consider the populationally identifiable sparse ANOVA decomposition:
$f_{0}(\mathbf{x})=\sum_{S\in\mathbb{S}}f_{0,S}(\mathbf{x}_{S}),$
where $f_{0,S},S\in\mathbb{S}$ satisfy the populational identifiability condition ($\mathbb{P}_{\mathbf{X}}$-identifiability condition) and $\mathbb{S}\subseteq\text{power}([p],d_{\max})$ is the index set of signal components. Let $d_{\max}=\max_{S\in\mathbb{S}}|S|$ that is the maximum order of signal components in the true function. We assume the following regularity conditions.
@@
    The density $p_{\mathbf{X}}$ of $\mathbb{P}_{\mathbb{X}}$ with respect to the Lebesgue measure on $\mathbb{R}^{p}$ exists and satisfies $0<\inf_{\mathbb{x}\in\mathcal{X}}p_{\mathbf{X}}(\mathbb{x})\leq\sup_{\mathbb{x}\in\mathcal{X}}p_{\mathbf{X}}(\mathbb{x})<\infty.$
@@
    Additionally, we assume that $\|f_{0,S}\|_{\infty}\leq F$ for some positive constant $F$. For these assumptions, we write $f_{0,S}\in\mathcal{H}^{\alpha}_{F}.$ Furthermore, we write $f_{0}\in\mathcal{H}^{\alpha}_{0,F}$ if $f_{0,S}\in\mathcal{H}^{\alpha}_{F}$ for all $S\in\mathbb{S}$.
@@
    The log-partition function $A(\cdot)$ is differentiable with a bounded second derivative over $\xi,\xi]$, i.e., there exists a positive constant $C_{A}$ such that
    $1/C_{A}\leq\ddot{A}(x)\leq C_{A}$
    for all $x\in[-\xi,\xi].$
@@
Theorem 5.1 (Posterior concentration rate of ANOVA-BART).
Assume that (J.1), (J.2),(J.3) and (J.4) hold. Then, for a given $\xi>2^{p}F,$ we have
$\pi_{\xi}\big\{f:\|f_{0}-f\|_{2,n}>B_{n}\epsilon_{n}\big|\mathbf{X}^{(n)},Y^{(n)}\big\}\xrightarrow{}0,$
for any $B_{n}\xrightarrow{}\infty$ in $\mathbb{P}_{0}^{n}$ as $n\xrightarrow{}\infty,$ where $\epsilon_{n}=n^{-\frac{\alpha}{2\alpha+d_{\max}}}(\log n)^{1\over 2}.$
Note that our prior is assigned to satisfy the $\mu_{n}$-identifiability condition, while the true function satisfies the $\mathbb{P}_{\mathbf{X}}$-identifiability condition. This discrepancy constitutes one of the technically delicate points of the proof of Theorem 5.1.
5.2 Comparison with Bayesian Tree Ensemble Models
Ročková and Saha (2019); Jeong and Ročlková (2023) derived the posterior concentration rates for BART, which are minimax optimal. Moreover, Ročková and van der Pas (2020) developed Bayesian forest, whose posterior concentration rate is adaptive to $d_{\max}$ as ANOVA-BART is, but its practical implementation would be computationally demanding.
ANOVA-BART adopts a fundamentally different strategy than BART. In particular, the number of decision trees increases as the sample size increases in ANOVA-BART while the sizes of each decision tree increase in BART. ANOVA-BART approximates $f_{0}$ using a sum of many small decision trees, whereas BART approximates $f_{0}$ using a finite number of large decision trees. That is, we have to prove that the sum of many small identifiable binary-product trees can approximate a smooth function well. For this purpose, we first approximate a smooth given function by a large decision tree as is done by Ročková and van der Pas (2020). Then, we prove that this large decision tree can be represented by the sum of many small identifiable binary-product trees, which is technically quite involved. See Section E of Supplementary Material.
An important practical advantage of ANOVA-BART compared to BART is Theorem 5.2 below, which derives the posterior concentration of each component. An obvious applications of Theorem 5.2 is to screen out unnecessary components a posteriori.
Theorem 5.2 (Posterior concentration rate of each component).
@@ 5.2 Comparison with Bayesian Tree Ensemble Models
Then, for $S\subseteq[p]$ and $\xi>2^{p}F,$ we have
@@
$\pi\{\|f_{S}\|_{2,n}>\epsilon_{n}\log n|\mathbf{X}^{(n)},Y^{(n)}\}<\delta$
for a given positive real number $\delta$. Theorem 5.2 implies that this post-hoc component selection procedure is selection-consistent (i.e. deleting all unnecessary components).
@@ 6 Experiments
This section presents the results of numerical experiments of ANOVA-BART. In Section 6.1, we conduct an analysis of synthetic data, while we focus on real data analysis in Section 6.2. In Section 6.3, we investigate the stability of ANOVA-BART in estimating the components.
We consider BART (Chipman et al. (2010)), SSANOVA (Gu and Wahba (1993)), MARS (Friedman (1991)) and NAM (Agarwal et al. (2021)) as the baseline methods. We use the official code in ‘BayesTree’ R package (Chipman et al. (2016)), ‘gss’ R package (Gu (2014)), ‘earth’ R package (Milborrow et al. (2017)) and the official code in https://github.com/AmrMKayid/nam for implementing BART, SSANOVA, MARS and NAM, respectively. For NAM, we extend the code to implement NAM to include the second order interactions. The detailed descriptions of the selection of hyperparameters in ANOVA-BART and baseline methods are presented in Section H of Supplementary Material. Note that while SSANOVA and NAM require specifying the maximum order of the components to be estimated, ANOVA-BART does not require the predefined maximum order. We set the maximum order of interaction at 2 for SSANOVA and NAM because a larger value than 2 would not be possible due to memory deficiency.
6.1 Analysis of synthetic data
We evaluate the performance of ANOVA-BART in view of prediction and component selection by analyzing synthetic data. To do this, we generate data using the Friedman’s test function (Friedman (1991); Chipman et al. (2010); Linero and Yang (2018)) defined as
@@ 6.1 Analysis of synthetic data
where $\epsilon\sim N(0,\sigma_{\epsilon}^{2})$. The error variance $\sigma_{\epsilon}^{2}$ is set to make the signal-to-noise ratio be 5. For training data, we generate $\mathbf{x}_{i}=(x_{i,1},...,x_{i,p})^{\top}$, i=1,,1000i=1,...,1000 from the uniform distribution on $(0,1)^{p}$ and then generate the response variable using only the first 5 covairates. This means that all covariates except the first five are noises.
@@ 6.1.1 Prediction performance
For all methods, hyperparameters are selected via 5-fold cross-validation on the training data. Independently generated test data of size 10,000 are used to compute the Root Mean Square Error (RMSE) as a measure of prediction performance. For ANOVA-BART and BART, we use the Bayes estimators.
Table 1 presents the averages and standard errors of the RMSEs for ANOVA-BART and baseline methods based on 5 repetitions of the simulation. The results amply support that ANOVA-BART is favorably compared to the baseline methods in prediction performance.
Table 1: Averages and standard errors of RMSEs on the synthetic data.
ANOVA-BART BART MARS SSANOVA NAM
p=10 1.188 (0.01) 1.174 (0.02) 1.196 (0.02) 1.207 (0.05) 1.207 (0.03)
p=50 1.224 (0.01) 1.256 (0.03) 1.310 (0.04) 1.550 (0.05) 1.713 (0.04)
p=100 1.275 (0.02) 1.304 (0.03) 1.407 (0.04) 1.702 (0.14) 1.779 (0.04)
@@ 6.1.2 Component selection
To examine how effectively ANOVA-BART selects the true signal components, we conduct a simulation. We use the $l_{2}$ norm of each estimated component (i.e., $\|f_{S}\|_{2,n}$) as the important score, i.e., if $\|f_{S}\|_{2,n}$ large, we consider $f_{S}$ to be important. We select the top 10 components based on the normalized important scores (the important scores divided by the maximum important score), which are presented in Figure 2 for various values of $p.$ The red bars correspond to the signal components, while the blue bars are noisy ones. The results strongly indicate that ANOVA-BART can detect signal components very well.
@@ 6.2 Real data analysis
Real data Size Dimension of covariates Task
Boston (Perera (2018)) 506 13 Regression
Abalone (Warwick et al. (1995)) 4,177 8 Regression
Servo (Ulrich (1986)) 167 4 Regression
Mpg (Quinlan (1993)) 392 7 Regression
Breast (Wolberg et al. (1993)) 569 30 Classification
Churn (59) 7,043 20 Classification
Madelon (Guyon (2004)) 2,600 500 Classification
We analyze 7 real data sets including Boston, Abalone, Servo, Mpg, Breast, Churn, and Madelon data sets. The four data sets Boston, Abalone, Servo and Mpg are analyzed in Chipman et al. (2010) while the remaining three data sets Breast, Churn and Madelon are examined in Park et al. (2025a). Table 2 summarizes the descriptions of the 5 data sets. We split each data into 80% training data and 20% test data, and repeat this random split 5 times to obtain 5 prediction measures.
@@ 6.2.1 Prediction performance
Table 3 presents the averages and standard errors of the RMSE and AUROC values for ANOVA-BART and baseline methods on 7 real data sets. Similar to the case with synthetic data, ANOVA-BART shows comparable prediction performance on real data as well. In particular, on the Madelon and Servo datasets, ANOVA-BART as well as BART and MARS outperform NAM and SSANOVA significantly. Note that the former three methods are able to detect higher order interactions while the later two methods are limited to estimating the second-order interactions. Table 4 presents the top 5 most important components according to the normalized importance scores defined in Section 6.1.2, which clearly shows that higher order signal interactions exist in the Madelon and Servo datasets. That is, NAM and SSANOVA fail to detect signal higher order interactions which result in inferior prediction performances. In contrast, ANOVA-BART detects such higher order interactions successfully. BART and MARS can detect higher-order interactions but BART is not easily interpretable and MARS is not good at uncertainty quantification.
Table 3: Averages and standard errors of RMSEs on real data sets.
Real data Measure ANOVA-BART BART MARS SSANOVA NAM
Boston RMSE $\downarrow$ 3.448 (0.59) 4.073 (0.67) 4.788 (0.67) 4.460 (0.65) 3.832 (0.67)
Abalone RMSE $\downarrow$ 2.112 (0.24) 2.197 (0.26) 2.137 (0.25) 2.137 (0.24) 2.062 (0.23)
Servo RMSE $\downarrow$ 0.316 (0.02) 0.342 (0.04) 0.441 (0.06) 0.820 (0.03) 0.802 (0.04)
Mpg RMSE $\downarrow$ 2.486 (0.32) 2.699 (0.43) 3.019 (0.52) 3.091 (0.47) 2.755 (0.41)
Breast AUROC $\uparrow$ 0.998 (0.001) 0.993 (0.002) 0.993 (0.003) 0.984 (0.004) 0.988 (0.001)
Churn AUROC $\uparrow$ 0.852 (0.007) 0.849 (0.006) 0.847 (0.007) 0.848 (0.007) 0.848 (0.008)
Madelon AUROC $\uparrow$ 0.866 (0.005) 0.765 (0.005) 0.863 (0.016) 0.553 (0.009) 0.644 (0.001)
Table 4: Top 5 important components.
Rank 1 Rank 2 Rank 3 Rank 4 Rank 5
Dataset Component Score Component Score Component Score Component Score Component Score
Servo {1} 1.000 {1,7,8} 0.461 {8} 0.434 {1,8} 0.324 {6} 0.298
Madelon {49, 319, 339} 1.000 {337} 0.951 {319, 454} 0.852 {106} 0.725 {49} 0.528
@@ 6.2.2 Uncertainty quantification
An important advantage of Bayesian methods compared to frequentist’s counterparts is superior performance of uncertainty quantification. We compare ANOVA-BART and BART in view of uncertainty quantification. As a measure of uncertainty quantification, for regression task, we consider Continuous Ranked Probability Score (CRPS, Gneiting and Raftery (2007)). For a given test sample $(\mathbb{x},y),$ CRPS is defined as
@@
where $F_{\mathbb{x}}$ is the predictive distribution of $Y$ given $\mathbb{X}=\mathbb{x}.$ For classification task, we consider Expected Calibration Error (ECE, Kumar et al. (2019)) as an uncertainty quantification measure. For evaluation, we typically report the average CRPS or ECE on the test data, where smaller values indicate better uncertainty quantification.
Table 5 presents the averages and standard errors of 5 CRPS and 5 ECE values obtained by 5 random splits of training and test data. Not surprisingly, the two Bayesian methods, BART and ANOVA-BART, show better performance than the frequentist method, MARS.
Table 5: Averages and standard errors of CRPS and ECE on real data sets.
Real data Measure ANOVA-BART BART MARS
Boston CRPS $\downarrow$ 2.368 (0.29) 2.623 (0.25) 3.580 (0.49)
abalone CRPS $\downarrow$ 1.351 (0.16) 1.384 (0.18) 1.575 (0.16)
servo CRPS $\downarrow$ 0.200 (0.02) 0.202 (0.02) 0.268 (0.03)
mpg CRPS $\downarrow$ 1.564 (0.17) 1.553 (0.27) 2.367 (0.47)
Breast ECE $\downarrow$ 0.066 (0.011) 0.071 (0.013) 0.072 (0.028)
Churn ECE $\downarrow$ 0.030 (0.002) 0.033 (0.003) 0.045 (0.001)
Madelon ECE $\downarrow$ 0.074 (0.004) 0.084 (0.008) 0.106 (0.005)
6.3 Stable interpretation of ANOVA-BART
Figure 3: The functional relations of the two estimated components by ANOVA-BART on Boston data.
Relations between the covariates and response variable can be interpreted by looking at the functional relations of each estimated component in ANOVA-BART. For illustration, Figure 3 presents the plots of the two estimated components of ANOVA-BART on Boston data. Specifically, each plot displays the Bayes estimate along with the 95% pointwise credible interval for ‘The average number of rooms in a dwelling’ and ‘The proportion of the lower-status population’. These plots suggest that an increase in ‘The average number of rooms in a dwelling’ is associated with a positive contribution to housing prices, whereas an increase in ‘The proportion of the lower-status population’ is associated with a negative contribution to housing prices.
For this interpretation to be useful, the estimated components should be stable in the sense that the result is robust to a small perturbation of traning data. In fact, ANOVA-BART employs the sum-to-zero condition to ensure that each component is identifiable and so can be estimated stably. In contrast, neural network based methods such as NAM ((Agarwal et al., 2021)) and NBM ((Radenovic et al., 2022)) are unstable in estimation of the component mainly due to the non-identifiability of each component.
Here, we conduct an experiment to compare the stability of ANOVA-BART in component estimation with that of NAM. For this experiment, we estimate the components based on a bootstrap dataset. This procedure is repeated five times to have five estimators for each component. We then compute the stability score using these estimators, where the stability score is defined as
$\mathcal{SC}^{d}(f)=\sum_{S\subseteq[p],|S|\leq d}\mathcal{SC}(f_{S}).$
Here, $\mathcal{SC}(f_{S})$ is defined as
$\mathcal{SC}(f_{S}):=\frac{1}{|\mathcal{C}_{S}|}\sum_{\mathbf{x}\in\mathcal{C}_{S}}\frac{\sum_{j=1}^{5}(\hat{f}_{S}^{j}(\mathbf{x})-\bar{f}_{S}(\mathbf{x}))^{2}}{\sum_{j=1}^{5}(\hat{f}_{S}^{j}(\mathbf{x}))^{2}},$
where $\hat{f}^{j}_{S}$ denotes the estimate of component $f_{S}$ obtained from the $j$th bootstrap data, $\bar{f}={1\over 5}\sum_{j=1}^{5}\hat{f}_{S}^{j}(\tilde{\mathbf{x}}_{i}),$ and $\mathcal{C}_{S}=\prod_{j\in S}\{x_{(0.09),j},x_{(0.19),j},\ldots,x_{(0.99),j}\}$ with $x_{j,(\alpha)}$ being the $\alpha\times 100\%$ quantile of the empirical distribution of $\{x_{1,j},\ldots,x_{n,j}\}.$ A lower stability score indicates a more stable estimation of the component.
Table 6 compares the stability scores $\mathcal{SC}^{2}(f)$ of ANOVA-BART and NAM for multiple benchmark datasets. These results confirm that ANOVA-BART outperforms the NAM in terms of the stability of component estimation, which suggests that enforcing the identifiability condition is important for stable interpretation.
Table 6: Stability scores of ANOVA-BART and NAM.
Real data ANOVA-BART NAM
Boston 0.435 0.705
abalone 0.452 0.770
servo 0.478 0.665
mpg 0.515 0.560
Breast 0.133 0.730
Churn 0.083 0.730